# Predicting Energy Budgets in Droplet Dynamics: A Recurrent Neural Network Approach

### v1

@@ C.1 Spreading dataset
To visualize the spreading behavior while varying Re and We, we have presented in Fig. 3 the droplet morphology and the temporal evolution of the diameter for shapes 1 and 10. As depicted in this figure, increasing the Reynolds number results in a greater spread for both shapes. This behavior is quantified by the subplots in Fig. 3 depicting the time variation of the droplet diameter. Figs. 3 and 3 illustrate snapshots comparing two shapes for two different sets of dimensionless numbers, while Figs. 3 and 3 compare the evolution of the spreading diameter through time for both shapes. It is evident from Figs. 3 and 3 that at lower Reynolds numbers, the droplets spread more slowly compared to the second test involving higher $Re$.
@@ C.2 Collision dataset
The second dataset consists of two drops of 1mm in diameter colliding horizontally with one another. The pair of drops merge after collision and generates horizontal and vertical oscillation over time, helping to create a time series that is harder to predict and that can be used to test the machine learning model’s efficiency.
To define the numerical simulation parameters for each sample, cases similar to those presented in the experimental coalescence map from [45] were selected by varying the impact velocity. For each set of Reynolds and Weber numbers, 20 simulations are run by varying the impact parameter B in the intervals shown in Table 3. Once again, certain samples were not successfully run due to solver errors. In total, 72 samples were generated.
@@
To visualize the difference in horizontal and vertical spreading after impact resulting from changes in the physical parameters and the impact parameter, snapshots from some simulated samples are shown in Fig. 4. Figs. 4, 4, and 4 illustrate snapshots of two coalescing droplets on three different sets of dimensionless numbers, with pairs 4 and 4 varying only the Reynolds and Weber numbers, and 4 and 4 varying only the impact parameter. For Figs. 4, 4, and 4, we compare the time series of diameter and height oscillation for each example.
@@ D.   Machine learning model
Since the transient features in the extracted data have a certain degree of time-correlation, it would be beneficial to use a neural network architecture that can capture it effectively. For this purpose, a Recurrent Neural Network [52], which can be used to find correlations on data sequences, is a good option.
The drawback of a simple RNN architecture is the likely introduction of the vanishing gradient problem [53], considering that the data has a large number of time steps. To avoid this, the LSTM [54], a type of RNN architecture employed to efficiently learn long-term dependencies in data, is chosen to be implemented.
@@
In contrast to the hidden state, which mostly carries information from the previous time step, the cell state is capable of passing data through multiples cells, thanks to its contents suffering only minor reductions and/or increments between each time step.
@@
where $\sigma$ is the sigmoid activation function, $W_{f}$ is a weight matrix, $[\bar{y}_{t-1},x_{t}]$ is the concatenation of the previous cell’s output with the current cell’s input and $b_{f}$ is a bias.
@@
The most interesting features to predict, whether in a numerical or in a experimental sense, are divided in two types depending on the output type, being static predictions, where each predicted feature has only one value, and transient predictions, where each predicted feature has its values changing through time. Our work focuses on the transient predictions, but an exploration of static predictions is also provided in Section C.2.
For the input, there can be both static and transient features. Since the input dimensions need to be the same for every feature, all of the static data are repeated so that each have multiple time steps with the same value. For example, for the input of the dimensionless number Re, the repetition would generate $Re(0)=Re(1)=\cdots=Re(t_{max})$.
For the hidden layers, we have a number of LSTM layers, followed by a Dense layer to generate values within the expected range for predictions and a Reshape layer to control the output shape. The schematics for the input/output relations of one of the explored neural network structures for the droplet collision problem is shown in Fig. 6.
@@
To better visualize the explored predictions, all test cases are shown in Tab. 4, characterized by output type (static or transient), dataset used, input features and output features.
@@
Type
Dataset
Input
Output
Transient
Spreading
$t,D(t)$, Re, We
$E_{k}(t),E_{s}(t),E_{d}(t)$
Transient
Collision
$D(t),H(t)$, Re, We, B
$E_{k}(t),E_{s}(t),E_{d}(t)$
Transient
Collision
$D(t),H(t)$
$E_{k}(t),E_{s}(t),E_{d}(t)$
Static
Collision
$E_{k}(t),E_{s}(t),E_{d}(t)$
Re, We
Table 4: Description of the explored test cases.
@@ A.   Performance of the model formulations
Firstly, as the hyperparameter tuning is a crucial stage for applying machine learning modeling, in the current work we have employed the open source hyperparameter optimization framework Optuna [55]. After a detailed investigation considering the datasets obtained by the simulations of droplets with varying initial shapes impacting on solid surfaces and collision of droplets, the hyperparameters were chosen as follows
@@
Despite this difference, both metrics still fall near the optimal ranges, even presenting higher accuracy with fewer training samples than the spreading dataset, as will be shown next. This also shows that predicting the behavior of a new drop shape (spreading case) is more complex than predicting the behavior of new flow parameters (collision case).
@@ B.   Prediction for the energy budget
To better visualize the model’s efficiency for the spreading case, we first analyse the prediction behavior of two different samples shown in Fig. 7. Each energy plot is normalized by the total energy in the first time step ($E_{t}(0)$).
In Figs. 7(a) and 7(b), we can observe pronounced fluctuations caused by the sliding technique (as depicted in Fig. 1) applied to the dataset. Given that each sample is divided into multiple sub-samples, the predictions can oscillate between each one of them, as observed in the surface energy $E_{s}$ shown in Fig. 7(a) and the viscous dissipation $E_{d}$ from Fig. 7(b). While it may appear to increase errors, some tests have shown that the technique, in fact, enhances model accuracy by generating more samples and using fewer time steps, which reduces training complexity.
@@
The predominance of samples within this lower range also influences the training process. Samples with lower Weber numbers, which have higher surface tension effects, exhibit higher surface energies compared to the majority. As depicted in Fig. 7(b), this higher surface energy undermines accuracy, as the model is primarily tuned for lower values. On the other hand, a higher Reynolds number can impact the prediction of kinetic energy $E_{k}$, affecting the capability of accurately predicting the curve in the few steps leading up to when the drop stops moving, as somewhat depicted in Fig. 7(b).
(a) 1st sample - Re = 30.73, We = 1659.38.
(b) 4th sample - Re = 24.81, We = 90.62.
@@
In general, the main difference in energy behavior observed in the spreading dataset lies in the steepness of its curves. Since the initial drop shape is one of the key factors influencing this behavior, the analysis indicates that the model is capable of generating accurate predictions for samples with a new drop shape (see Fig. 2).
For the collision dataset, we illustrate in Fig. 8 the prediction behavior of two different test cases. Again, the energy plot is normalized using the total energy in the first time step.
The collision adds another layer of complexity to the problem, as the energies now exhibit oscillations over time. In Fig. 8(a) we observe that the model can produce accurate predictions even when there is an anomaly in the curves (as seen at time $t\approx 6$). Furthermore, although fluctuations in viscous dissipation $E_{d}$ are still detectable, they occur much less frequently. Decreasing Re and We, as illustrated in Fig. 8(b), the coalesced droplet experiences an oscillatory regime that can also be correctly captured by the prediction. Overall, according to these figures, we can confirm the excellent agreement between the expected and predicted results.
(a) 1st sample - Re = 115.56, We = 53.0, B = 0.17.
(b) 21st sample - Re = 74.45, We = 22.0, B = 0.27.
Fig. 8: Comparison between the expected and predicted energy time series from three samples of the collision test dataset.
@@ C.   Two-phase prediction with sequential neural networks
To further investigate the significance of our proposed methodology, we adopt a two-phase prediction strategy tailored specifically for the droplet collision problem. Initially, we consider only the drop diameter $D(t)$ and height $H(t)$ as inputs to our model. Through this initial phase, our aim is to accurately estimate the energy budget associated with the collision process. Subsequently, leveraging the insights gained from the energy budget prediction, we extend our model to predict static values, such as the nondimensional parameters Re and We.
This two-phase prediction approach holds particular relevance in experimental settings where obtaining precise fluid parameters is challenging. Often, despite having access to video data from experiments, the lack of comprehensive knowledge regarding the fluid properties impedes accurate prediction of the energy budget. In such scenarios, our method offers a robust solution by including geometric data to not only predict the energy budget but also estimate essential dimensionless parameters crucial for understanding the underlying physics of droplet dynamics. Using this innovative two-phase prediction process, we bridge the gap between experimental observations and theoretical insights, thereby enhancing our ability to analyze and interpret droplet collision phenomena with greater precision and efficiency.
@@ C.1 Energy budget prediction
Although the model experiences a decrease in accuracy when the dimensionless numbers are removed, the overall behavior of the energies is still correctly captured. Moreover, given the extremely limited availability of collectable data from just a video, this increase in error remains tolerable. To assess the differences in metrics across training and test samples for both configurations, let us examine Table 6.
@@ C.2 Dimensionless numbers prediction
In a parallel test, we attempted to predict the dimensionless numbers directly using only geometric data. However, this approach resulted in poor prediction accuracy for the collision dataset, as shown in Appendix B.
Therefore, rather than directly using $D(t)$ and $H(t)$ to predict the dimensionless numbers Re and We, the second step involves utilizing the previously trained neural network to predict the energies and then training a new neural network to predict the dimensionless numbers using these energies.
Given that each sample is divided into 100-time-step slices, the selected prediction for a sample is determined by the median of the predicted values for each of its 10 slices. This is done to mitigate potential noise caused by high errors resulting from using approximated energies as inputs. The prediction results for the collision test dataset are depicted in Fig. 11.
@@
Metric
Training samples
Test samples
$R^{2}$
0.99366
0.99823
$RMSE$
2.50602
1.39455
$NRMSE$
0.02534
0.0141
@@
With this, we have demonstrated that, in certain regimes, it may be feasible to predict energies and dimensionless numbers using only geometric data extracted from experiment videos. As a follow-up to this study, new architectures could be developed and tested with additional datasets to enhance the neural network generalization, thereby expanding its applicability to a wider range of regimes.
@@ IV.   Conclusions
Initially, we proposed two LSTM models to predict the energy balance in two moving free surface problems: the impact of non-spherical droplet in a rigid surface and the collision of two equal-sized droplets. Since the first dataset required generating data from different and unusual shapes, a method capable of extracting particles from interfaces present in digital images was also developed (as shown in Appendix A). Following this test, a two-phase sequential prediction approach was applied in an attempt to predict important properties using exclusively geometric data. This method is specially useful for experimental observations, since it only uses data which can be easily extracted from experiment images or videos. It provides an easy-to-use tool for the challenging task of predicting fluid parameters such as the energies and the dimensionless numbers.
@@
In the context of one-phase prediction, both models shared the same overall architecture, with the only variation being in the data used for training. The first model was trained using data from 95 numerical simulations involving seven different shapes, extracted from images of real experiments, impacting a solid surface and spreading over time. The second model utilized data from 50 numerical simulations of two colliding drops that spread and retracted after coalescence, causing oscillations in the energy budget. For the droplet impact problem, the inputs to the neural network consist of simulation time, the temporal series of the droplet diameter and nondimensional numbers. However, for the droplet collision scenario which is more complex than the spreading case, the model is initialized with the temporal series of diameter and height of the coalesced droplet, in combination with nondimensional numbers and the impact parameter. In both models, the output corresponds to the energy budget.
@@
To evaluate the performance of each model, we conducted a comprehensive metric analysis commonly employed in machine learning modeling, which includes the Coefficient of Determination ($R^{2}$), the Root Mean Square Error ($RMSE$), and the Normalized Root Mean Square Error ($NRMSE$). The $R^{2}$, $RMSE$ and $NRMSE$ metrics for both models are, respectively, 0.99429 and 0.99559, 0.02957 and 0.0137, and 0.02961 and 0.02096. After analyzing the metrics, we can observe that both models achieved performances sufficiently close to the ideal values for each metric, with $R^{2}s$ close to one and $RMSEs$ approaching zero. The analyzed models are accurate in a similar way, even when applied to considerably different problems, which can be evaluated by the similarity of their $NRMSEs$. The spreading model showed promising results by successfully handling a completely new shape. Similarly, the collision dataset’s model performed well with samples featuring sets of parameters not used during training. The accuracy of the models is also evaluated through plots depicting the kinetic, surface, and viscous dissipation, confirming excellent agreement between the simulation data and the predicted values. This suggests that the machine learning architectures have the potential to be highly adaptable and accurate across a range of fluid regimes.
@@
According to the results for a two-phase prediction, the architecture remains similar, but now follows a two-step scheme. Firstly, an LSTM network predicts the energy budget using only geometric data. Then, the predicted energies serve as inputs in a second LSTM network to predict the static dimensionless numbers. The metrics for the first phase (energies) and the second phase (static prediction for Reynolds and Weber numbers) are, respectively, 0.9776 and 0.99823, 0.03087 and 1.39455, and 0.04725 and 0.0141. Despite a slight increase in error when compared to the one-phase prediction strategy, the metrics for two-phase predictions still demonstrate high accuracy. After the performance investigation concerning the metrics, we have also included the results for predicting the Reynolds and Weber numbers. Results show excellent agreement for low values of Re and We, whereas for larger values, the neural network appears to have more difficulty in establishing accurate predictions. This difficulty can be attributed to the fact that, for the training data, there are fewer instances with higher Re and We values, as seen in Table 3. Finally, this approach proves useful in the sense that directly attempting to predict the dimensionless numbers yielded higher errors, as demonstrated in Appendix B.
While the metrics indicate high versatility, there are still some questions which must be explored. One notable task is the challenge of predicting regimes beyond coalescence in the collision model, given that our current model predominantly relies on coalescence data. For instance, capturing the distinct behaviors associated with the bouncing regime poses a significant challenge, as transitioning between regimes presents a complex task for machine learning models. Future research could be focused on finding new appropriate fluid parameters specific to the bouncing phenomena (or any other regime, for that matter) and exploring interpretation and generalization of the model.
@@
We hope that this research will shed light on the advantages of leveraging the performance offered by LSTM architectures for predicting different types of transient data in complex flows of complex fluids, as for instance dynamics of viscoelastic droplets [56, 57].

### v2

@@ C.1 Droplet spreading dataset
To visualize the spreading behavior while varying Re and We, Fig. 3 illustrates the droplet morphology and the temporal evolution of the diameter for shapes 1 and 10. As depicted in this figure, increasing the Reynolds number results in a greater spread for both shapes. This behavior is quantified by the subplots in Fig. 3 depicting the time variation of the droplet diameter. Figs. 3 and 3 illustrate snapshots comparing two shapes for two different sets of dimensionless numbers, while Figs. 3 and 3 compare the evolution of the spreading diameter through time for both shapes. Figs. 3 and 3 demonstrate that at lower Reynolds numbers, the droplets spread more slowly compared to the second test involving higher $Re$.
@@ C.2 Collision dataset
The second dataset consists of two drops of 1mm in diameter colliding horizontally. After the collision, the droplets merge and generate both horizontal and vertical oscillations over time, creating time series that are more challenging to predict and that can be used to test the efficiency of the machine learning model.
To define the numerical simulation parameters for each sample, cases similar to those presented in the experimental coalescence map from [Karrar2019] were selected by varying the relative impact velocity, with the characteristic length $L$ defined as the initial diameter of one of the droplets. For each set of Reynolds and Weber numbers, 20 simulations are run by varying the impact parameter B in the intervals shown in Table 3. In total, 72 simulations were performed, following the same setup as before, with each running until $t_{\text{max}}=50$ and producing 1000 equally spaced snapshots.
@@
The effect of parameters on horizontal and vertical elongations after impact is depicted through simulation snapshots in Fig. 4. Figs. 4, 4, and 4 illustrate snapshots of two coalescing droplets on three different sets of dimensionless numbers, with pairs 4 and 4 varying only the Reynolds and Weber numbers, and 4 and 4 varying only the impact parameter. For Figs. 4, 4, and 4, we compare the time series of diameter and height oscillation for each example.
@@ D.   Machine learning model
In this work, we propose a machine learning model based on a recurrent neural network to predict the temporal variation of the energy budget. Given the complex nature and the underlying temporal dependency of data in this work, we have adopted the Long Short-Term Memory (LSTM) model[Hochreiter1997]. With a geometric input data set, the model learns a mapping to produce temporal predictions in terms of the energy budget. The term “geometric” refers to the time variation of the droplet diameter during spreading on a surface or the diameter and height of the coalesced droplet after the collision of two droplets. Incorporating dimensionless parameters such as the Reynolds and Weber numbers, as well as geometrical quantities as the collision impact parameter, enhances the accuracy of predictions for the transient behavior of energy components.
Since the transient features in the extracted data have a certain degree of time-correlation, it would be beneficial to employ a neural network architecture that can capture it effectively. For this purpose, a Recurrent Neural Network [Jain1999], which is effective at finding correlations on data sequences, was considered as an option.
The drawback of a simple RNN architecture is the likely introduction of the vanishing gradient problem [Seol2021], considering that our data has a large number of time steps. To address this issue, we selected the LSTM architecture [Ralf2019], a variant of RNN designed to efficiently learn long-term dependencies in data.
@@
Unlike the hidden state, which mostly carries information from the previous time step, the cell state is capable of passing data through multiples cells with minimal reductions and/or increments between time steps.
@@
where $\sigma$ is the sigmoid activation function, $W_{f}$ is a weight matrix, $[\bar{y}_{t-1},x_{t}]$ is the concatenation of the output from the previous cell with the input of the current cell and $b_{f}$ is a bias.
@@
The most interesting features to predict, whether in a numerical or in a experimental scenario, are divided in two categories depending on the output type: (i) static predictions, where each predicted feature has only one value, and (ii) transient predictions, where each predicted feature has its values changing through time. Our work focuses on the transient predictions, but an exploration of a static case is also provided in Section C.2.
Both static and transient features can be used in the network input layer. Since the input dimensions need to be the same for all features, all static data is transformed into constant time series over all time steps. For example, when the dimensionless number $Re$ is used as an input, it is converted into a constant time series $Re(0)=Re(1)=\cdots=Re(n_{max})=Re$.
The architecture consists of LSTM hidden layers, followed by a Dense layer to generate values within the expected range for predictions, and a Reshape layer to adjust the output shape. The schematics for the input/output relations of one of the explored neural network structures for the droplet collision problem is shown in Fig. 6.
@@
Table 4 summarizes the test cases, characterized by output type (static or transient), dataset used, input features and output features.
@@
Type Dataset Input Output
Transient Spreading $t,D(t)$, Re, We $E_{k}(t),E_{s}(t),E_{d}(t)$
Transient Collision $D(t),H(t)$, Re, We, B $E_{k}(t),E_{s}(t),E_{d}(t)$
Transient Collision $D(t),H(t)$ $E_{k}(t),E_{s}(t),E_{d}(t)$
Static Collision $E_{k}(t),E_{s}(t),E_{d}(t)$ Re, We
Table 4: Summary of the explored test cases.
@@ A.   Performance of the model formulations
Firstly, as the hyperparameter tuning is a crucial stage for applying machine learning modeling, in the current work we have employed the open source hyperparameter optimization framework Optuna [Optuna2024]. Following a detailed analysis of datasets from simulations of droplets with varying initial shapes impacting on solid surfaces and droplet collisions, the hyperparameters were chosen as follows
@@
Despite these differences, both metrics remain near optimal ranges, even presenting higher accuracy with fewer training samples than the spreading dataset, as will be shown next.
@@
To better emphasize the need for a more complex model such as the LSTM, we compared predictions generated by linear regression, a simple RNN, and our LSTM on the collision dataset.
As shown in the metrics in Table 6 and the prediction example in Fig. 7, the LSTM model (Fig. 7(c)) is essential for accurately capturing the complex behavior of the analyzed dynamics. While linear regression (Fig. 7(a)) fails to capture the temporal dynamics of this problem, the RNN (Fig. 7(b)) partially succeeds but still struggles to maintain context over many time steps.
Metrics Linear regression RNN LSTM
$R^{2}$ 0.9201 0.9652 0.99559
$RMSE$ 0.0583 0.0385 0.0137
$NRMSE$ 0.0892 0.0589 0.02096
Table 6: Comparison of the evaluation metrics for each technique.
(a) Linear regression
(b) RNN
(c) LSTM
Fig. 7: Comparison of the predictions given by different machine learning models. All cases correspond to a binary droplet collision with Re = 115.56, We = 53.0, B = 0.17.
Finally, another important aspect is the computational feasibility of the method. To assess the difference in computational cost between performing a full numerical simulation up to $t_{\text{max}}=50$ for each parameter and obtaining energy predictions through a trained model, time evaluation tests were conducted on similar hardware. Each numerical simulation, for both the spreading and collision cases, requires approximately 8 to 12 hours, depending on the complexity of the simulated dynamics. The offline training phase of all models takes 3 to 6 hours, depending on the early stopping tolerance, while the online testing phase takes, on average, 0.0729s to predict all 1000 time steps up to $t_{\text{max}}=50$ for each parameter set.
@@ B.   Prediction for the energy budget
To better visualize the model’s efficiency for the spreading case, we first analyze the prediction behavior of two different samples shown in Fig. 8. Each energy plot is normalized by the total energy in the first time step $E_{t}(0)$.
In Figs. 8(a) and 8(b), we can observe pronounced fluctuations caused by the sliding technique (as depicted in Fig. 1) applied to the dataset. Given that each sample is divided into multiple sub-samples, the predictions can oscillate between each one of them, as observed in the surface energy $E_{s}$ shown in Fig. 8(a) and the viscous dissipation $E_{d}$ from Fig. 8(b). While it may appear to increase errors, some tests have shown that the technique, in fact, enhances model accuracy by generating more samples and using fewer time steps, which reduces training complexity.
@@
The predominance of samples within this lower range also influences the training process. Samples with lower Weber numbers, which have higher surface tension effects, exhibit higher surface energies compared to the majority. As depicted in Fig. 8(b), this higher surface energy undermines accuracy, as the model is primarily tuned for lower values.
(a) Re = 30.73, We = 1659.38.
(b) Re = 24.81, We = 90.62.
@@
In general, the main difference in energy behavior between different entries in the spreading dataset lies in the steepness of their curves (see Fig. 2). Since the initial drop shape is one of the key factors influencing this behavior, the analysis in 8 indicates that the model is capable of generating accurate predictions for samples with a new drop shape.
For the binary collision dataset, we illustrate in Fig. 9 the prediction behavior of three different test cases. Again, the energy plot is normalized using the total energy in the first time step.
The collision adds another layer of complexity to the problem, as the energies now exhibit oscillations over time. In Fig. 9(a) we observe that the model can produce accurate predictions even when there is an anomaly in the curves (as seen at time $t\approx 6$). Furthermore, although fluctuations in viscous dissipation $E_{d}$ are still detectable, they occur much less frequently. Decreasing Re and We, as illustrated in Fig. 9(b), the coalesced droplet experiences an oscillatory regime that can also be correctly captured by the prediction. Overall, as shown in these figures, we can confirm the excellent agreement between the expected and predicted results in the interpolation scenarios.
All samples corresponding to $Re=120.45$ were reserved for the test set to evaluate the model’s performance in extrapolation scenarios. Fig. 9(c) demonstrates that the model still provides a reasonable qualitative prediction for the extrapolated value.
(a) Re = 115.56, We = 53.0, B = 0.17.
(b) Re = 74.45, We = 22.0, B = 0.27.
(c) Re = 120.45, We = 58.0, B = 0.17.
Fig. 9: Comparison between the expected and predicted energy time series from three samples of the collision test dataset. Subfigures (a) and (b) correspond to interpolation cases, with Re values within the training range of [74.45, 115.56], while subfigure (c) represents an extrapolation case with Re = 120.45.
To ensure the robustness of our machine learning strategy in addressing a physical constraint, we have included results concerning the total energy conservation. To illustrate this, Fig. 10 presents a comparison between the predicted and expected behavior of the total energy for different parameters. Due to numerical errors in the training data, which arise from using an intermediate mesh refinement, some slight fluctuations in energy over time can be observed. As the learning algorithm attempts to model these dynamics, the generated forecasts tend to replicate this fluctuating pattern, but mostly maintain the total energy. It is important to highlight that according to our tests, no physically impossible cases were observed in the predictions.
(a) Re = 74.45, We = 22.0, B = 0.19.
(b) Re = 85.48, We = 29.0, B = 0.19.
Fig. 10: Examples of total energy behavior for different parameter sets.
@@ C.   Two-phase prediction with sequential neural networks
To further investigate the versatility of our proposed methodology, we adopt a two-phase prediction strategy tailored specifically for the droplet collision problem. Initially, we consider only the droplet diameter $D(t)$ and height $H(t)$ as inputs to our model. In this initial phase, our aim is to accurately estimate the energy budget associated with the collision process. Subsequently, leveraging the insights gained from the energy budget prediction, we extend our model to predict static values, such as the nondimensional parameters $Re$ and $We$.
This two-phase prediction approach is particularly valuable in experimental contexts where precise fluid parameters measurements are difficult to obtain. Often, despite having access to video data from experiments, the lack of comprehensive knowledge regarding the fluid properties impedes accurate prediction of the energy budget. In such scenarios, our method provides a robust solution by including geometric data to not only predict the energy budget but also estimate dimensionless parameters essential for understanding the underlying physics of droplet dynamics. Using this innovative two-phase prediction process, we bridge the gap between experimental observations and theoretical insights, thereby enhancing our ability to analyze and interpret droplet collision phenomena with greater precision and efficiency.
@@ C.1 Energy budget prediction
While omitting the dimensionless numbers reduces the model’s accuracy, it still effectively captures the overall behavior of the energies. Moreover, given the high difficulty of measuring energy data from experimental videos through traditional methods, this increase in error remains tolerable. The differences in metrics across training and test samples for both configurations can be seen in Table 7.
@@ C.2 Dimensionless numbers prediction
In a parallel test, we attempted to predict the dimensionless numbers directly using only geometric data. However, this approach resulted in poor prediction accuracy for the collision dataset, likely due to the weak correlation between these terms, as discussed in Appendix B.
Therefore, rather than directly using $D(t)$ and $H(t)$ to predict the dimensionless numbers $Re$ and $We$, the second step involves first utilizing the previously trained neural network to predict the energies, then training a new neural network to predict the dimensionless numbers based on these energies.
Given that each sample is divided into 100 time step slices, the selected prediction for a complete sample (consisting of 1000 time steps) is determined by taking the median of the predicted values from its 10 slices. This is done to mitigate potential noise caused by high errors resulting from using approximated energies as inputs. The prediction results for the collision test dataset are depicted in Fig. 13.
@@
Metric Training samples Test samples
$R^{2}$ 0.99366 0.99823
$RMSE$ 2.50602 1.39455
$NRMSE$ 0.02534 0.0141
@@
With this, we have demonstrated that, in the explored regimes, it is feasible to predict energies and dimensionless numbers using only geometric data. This input data is easily accessible from experimental videos, so that this approach can also be used in experiments. As a follow-up to this study, new architectures could be developed and tested with additional datasets to enhance the neural network generalization, thereby expanding its applicability to a wider range of regimes.
@@ IV.   Conclusions
In this study, we proposed an LSTM-based model to predict the energy balance in two droplet dynamics problems: the impact of a non-spherical droplet on a rigid surface and the collision of two equal-sized droplets. For the first case, we also implemented a method to extract a Front-Tracking interface from experimental images allowing the use of realistic non-spherical droplet shapes in generating training data (see Appendix A). Following this test, a two-phase sequential prediction approach was applied to predict important properties using exclusively geometric data. The current methodology provides an easy-to-use tool for the challenging task of predicting important flow quantities such as the energy budget and the dimensionless numbers, which are useful for various industrial applications. As an example, optimizing atomizers or sprays for coating applications involves predicting whether the kinetic energy will be sufficient for the droplets to cover the entire surface and spread uniformly, and controlling the final droplet shape after impact, determining whether it will spread or maintain a more spherical form, through surface energy [Nasser2011]. Additionally, in agriculture, analyzing surface energy as a metric to understand droplet interactions with surfaces, such as their behavior on hydrophobic leaves, could help improve the control of agricultural sprayers, while evaluating viscous dissipation to assess energy loss of droplets upon impact may also contribute to optimizing sprayer design [Athira2024]. Finally, the food industry can also be mentioned, with the goal of modeling droplet behavior and controlling phenomena such as emulsion formation and droplet stability [Andrade2013].
@@
For the one-phase predictions, both proposed models share the same overall architecture, differing only in the data used for training. The first model was trained using data from 95 numerical simulations with seven different shapes extracted from real experimental images of droplets impacting on solid surfaces. In this model, the neural network input consists of simulation time, the temporal series of the droplet diameter and nondimensional numbers. The second model utilized data from 50 numerical simulations of two colliding droplets that elongate and retract after coalescence, leading to oscillations in the energy budget. Due to the more complex oscillatory behavior, this model is initialized with the temporal series of diameter and height of the coalesced droplet, in combination with nondimensional numbers and the impact parameter. In both models, the output is the energy budget over time.
@@
To evaluate the performance of each model, we conducted a comprehensive metric analysis commonly employed in machine learning modeling, which includes the Coefficient of Determination ($R^{2}$), the Root Mean Square Error ($RMSE$), and the Normalized Root Mean Square Error ($NRMSE$). The $R^{2}$, $RMSE$ and $NRMSE$ metrics for both models are, respectively, 0.99429 and 0.99559, 0.02957 and 0.0137, and 0.02961 and 0.02096. After analyzing the metrics, we observe that both models achieved performances sufficiently close to the ideal values for each metric, with $R^{2}$ close to one and $RMSE$ approaching zero. The analyzed models are accurate in a similar way, even when applied to considerably different problems, which can be evaluated by the similarity of their $NRMSEs$. The spreading model showed promising results by successfully handling a completely new shape. Similarly, the model for binary droplet collision performed well with samples featuring sets of parameters not used during training. The accuracy of the models is also evaluated through plots depicting the kinetic, surface, and viscous dissipation, confirming excellent agreement between the simulation data and the predicted values. This suggests that the machine learning architectures have the potential to be highly adaptable and accurate across a range of fluid regimes.
@@
For the prediction of nondimensional numbers, the model architecture remained similar but followed a two-phase scheme. Firstly, an LSTM network predicts the energy budget using only geometric data. Then, the predicted energies serve as inputs in a second LSTM network to predict the static dimensionless numbers. The metrics for the first phase (energies) and the second phase (static prediction for Reynolds and Weber numbers) are, respectively, 0.9776 and 0.99823, 0.03087 and 1.39455, and 0.04725 and 0.0141. Despite a slight increase in error when compared to the one-phase prediction strategy, the metrics for two-phase predictions still demonstrate high accuracy. Results show excellent agreement for low values of Re and We, whereas the neural network has more difficulty in predicting high values of $Re$ and $We$. This difficulty can be attributed to the fact that, for the training data, there are fewer instances with higher Re and We values, as seen in Table 3. Finally, this approach proves useful in the sense that directly attempting to predict the dimensionless numbers yielded higher errors, as demonstrated in Appendix B.
While the metrics indicate high versatility, some questions must still be explored. One notable challenge is the prediction of regimes beyond coalescence in the binary droplet collision scenario, as transitioning between regimes presents a complex task for machine learning models. This difficulty is further accentuated in our model due to its reliance on geometric data ($D(t)$ and $H(t)$) as inputs, which are significantly affected by regimes with bouncing or breakup. Future research could focus on the interpretation of the model itself, as well as modifications to the LSTM architecture, such as the Bidirectional LSTM [Sima2019], Attention LSTM [Xianyun2023], or adaptations to make it physics-informed [Dong2024].
@@
We hope that this research will shed light on the advantages of leveraging the performance offered by LSTM architectures for predicting different types of transient data in complex flows. Although our current work does not address complex fluids, it establishes a foundation for future studies that could extend these methods to address more challenging scenarios, such as the dynamics of viscoelastic droplets [Oishi2019, Hugo2022_2].